OpenStudy (sparklestaraa):

two questions as review from last semester i had question about - How to write the domain of a function? How to write the range of a function?

OpenStudy (amorfide):

domain of f is all the possible x values the function can have input the range is all the possible values that are output from the function

OpenStudy (sparklestaraa):

ok so like in equation form how might this look?

OpenStudy (sparklestaraa):

@hysusonic

OpenStudy (misty1212):

this would be lots easier if you had an example

OpenStudy (misty1212):

if the function is written as an equation, then you would write the domain probably in interval notation if it is given by ordered pair, then it would be a list

OpenStudy (amorfide):

depends on the function if you have function f(x)=x+2 you know all x values can be from negative infinite to infinite so your domain is all x that exist in reals domain of f:\[(-\infty, \infty)\] or \[\forall x \epsilon \mathbb{R} \]

OpenStudy (sparklestaraa):

ohhh so domain is like an ordered pair i think i understand now :) thanks :D I'm going to try to review with worksheets online :P When I'm done can someone maybe check? I am in semester two but i found this to be my hardest topic in semester 1 thats why I'm trying to get some clarification :P lol

OpenStudy (misty1212):

no dear the domain is not an ordered pair

OpenStudy (misty1212):

the domain is an INTERVAL usually

OpenStudy (misty1212):

you can write an interval as say \(0<x<10\) or \((0,10)\) but that means the interval, not the point

OpenStudy (sparklestaraa):

ohhhh give me an example of a function and i will try to make it a domain @misty1212

OpenStudy (misty1212):

like a function given by an equation? like say \[f(x)=\frac{x+1}{x-2}\]?

OpenStudy (sparklestaraa):

thanks misty xD thanks so much xD thats hard lol ok let me think whats the first number of the function like f(0) like 0 = what?

OpenStudy (misty1212):

for the one i wrote above, you cannot divide by zero, so you would set \(x-2=0\) and solve for \(x\) then the domain would be all numbers except that one

OpenStudy (sparklestaraa):

f(3) so all real numbers except 3?